Basic Scientific Computing

Marek Rychlik August 24, 2003

1 Welcome to Scientific Computing

Scientific computing papers have lots of formulas, like this one:

∑_{n=0}^{∞} \frac{1}{n} = ∞.

Mathematical formulas is one thing, tables are another. This is an example how we can format some data:

Course  Enrollment
math 481 16
math 581 2

2 Time for the next section

This section will have subsections.

2.1 The first subsection

Let us cite some scientific works. This is a well-known book on numerical analysis: [1].

References

[1] R. L. Burden and J. Faires. Numerical Analysis. Prindle, Weber & Schmidt, Boston, 4th edition, 1997.